# Chronos: Learning the Language of Time Series

### v1

@@ 1 Introduction
For the development of a useful general-purpose time series forecasting model, the scarcity of publicly available time series datasets, both in quantity and quality, is arguably more critical than the modeling framework. In addition to the comprehensive collection of public datasets we used to train Chronos, a central aspect of our approach is the integration of data augmentation strategies, including TSMix and KernelSynth. TSMix randomly samples a set of base time series from different training datasets, and generates new time series based on a convex combination of them; KernelSynth uses Gaussian processes to generate synthetic time series by randomly composing kernel functions. These techniques address the inherent limitations of small training datasets in time series forecasting, enhancing model robustness and generalization.
@@ 2 Background and Related Work
Time series forecasting can be addressed with a variety of different methods which can be broadly categorized into classical forecasting methods and deep learning methods. Classical forecasting methods such as ETS, ARIMA (Hyndman et al., 2008), Theta (Assimakopoulos & Nikolopoulos, 2000) fit a separate model to each time series independently (hence referred to as local models). In contrast, deep learning forecasting models learn across time series in a given dataset (and are called global models). These methods leverage advances in deep learning, such as RNNs which are used by DeepState (Rangapuram et al., 2018), DeepAR (Salinas et al., 2020), TimeGrad (Rasul et al., 2021), and transformers which are used by TFT (Lim et al., 2021) and PatchTST (Nie et al., 2023). Apart from the choice of architecture, these approaches differ in the way they model the target, with some modeling the density function while others directly predicting a set of quantiles (Wen et al., 2017; Gasthaus et al., 2019). Nevertheless, not all models produce probabilistic forecasts: notably, models such as Informer (Zhou et al., 2021) and DLinear (Zeng et al., 2023) only produce point forecasts.
@@ 3.2 Objective Function
Opting for a categorical output distribution offers two key advantages. Firstly, it requires no modification to the language model architecture or training objective, enabling the use of popular language modeling libraries and the utilities they provide out of the box (Wolf et al., 2020). Secondly, it imposes no restrictions on the structure of the output distribution, allowing the model to learn arbitrary distributions including multimodal ones. This flexibility proves especially valuable for a pretrained model, as time series datasets from diverse domains may follow distinct output distribution patterns.
@@ 4 Data Augmentation
4.1 TSMix: Time Series Mixup
Figure 2: An illustration of TSMix augmentation for $k=\{1,2,3\}$. TSMix improves pattern diversity by taking weighted combinations of randomly-sampled time series from different datasets.
Mixup (Zhang et al., 2017) is a data augmentation scheme proposed in the context of image classification. It generates convex combinations of random image pairs and their labels from the training dataset, which alleviates issues such as memorization and overfitting in deep learning models. Existing works (Carmona et al., 2021; Zhou et al., 2023b) have extended Mixup to the time series domain. Building upon these works, we propose TSMix, which generalizes the idea of Mixup to more than two datapoints. Concretely, TSMix randomly samples $k\sim\mathcal{U}\{1,K\}$ time series of a specific length, $l\sim\mathcal{U}\{l_{\min},l_{\max}\}$, from the training datasets, scales them, and takes their convex combination,
$\tilde{{\bm{x}}}^{\mathrm{TSMix}}_{1:l}=\sum_{i=1}^{k}\lambda_{i}\tilde{{\bm{x}}}^{(i)}_{1:l},$ (3)
where $\tilde{{\bm{x}}}^{(i)}_{1:l}$ denotes the $i$-th scaled time series. The combination weights, $[\lambda_{1},\dots,\lambda_{k}]$, are sampled from a symmetric Dirichlet distribution, $\mathrm{Dir}(\alpha)$. The complete pseudocode of TSMix can be found in Algorithm 1 in Appendix A. Intuitively, TSMix enhances the diversity of data by combining patterns from different time series. Figure 2 shows example augmentations generated by TSMix and illustrates how different patterns are mixed.
@@ 4.2 KernelSynth: Synthetic Data Generation using Gaussian Processes
While TSMix improves pattern diversity, it may still prove insufficient for training a generalist time series model, especially when real data is limited. To further supplement the training dataset, we propose KernelSynth, a method to generate synthetic time series using Gaussian processes (GPs). KernelSynth is inspired by the Automatic Statistician (Duvenaud et al., 2013), where a compositional search over a space of GP kernels is performed to explain the structure of a time series. We use the inverse of this process — randomly compose GP kernels to generate new time series.
@@ 5 Experiments
Pretraining-only
13
795,936
pretraining
Benchmark I
15
97,272
pretraining and in-domain evaluation
Naive, SeasonalNaive, AutoETS, AutoARIMA, AutoTheta, DeepAR, TFT, PatchTST, DLinear, WaveNet, N-BEATS, N-HiTS, GPT4TS
Benchmark II
27
103,047
zero-shot evaluation
All the above and ForecastPFN
@@ 5.2 Training Corpus and Protocols
We trained T5 models of 4 sizes, [^1] Our inference code and model checkpoints are available at https://github.com/amazon-science/chronos-forecasting. namely, Mini (20M), Small (46M), Base (200M) and Large (710M), and the GPT-2 base model (90M), on 10M TSMix augmentations (see Section 4.1) generated from the 28 training datasets, with $K=3$ in Algorithm 1, and 1M synthetic time series generated using Gaussian processes (see Section 4.2). Note that with this setup, original time series are adequately represented since they are included in the TSMix augmentations with probability $1/3$. We sampled time series from the augmentations and synthetic data in the ratio 9:1 during training. Each model is trained with an effective batch size of 256 sequences, using distributed data parallelism and gradient accumulation, whenever necessary. These sequences are constructed by slicing random windows from the time series, and then scaling and quantizing them into equal-sized bins within the interval $[l{=}-15$, $r{=}15]$, as described in Section 3.1. The context length of the sequences was set to 512, the default for T5 models, and the prediction length is set to 64, a value greater than the prediction lengths of all tasks we consider in our evaluation.
@@ 5.3 Baselines
We assessed the performance of Chronos models against a variety of time series forecasting baselines. From statistical forecasting literature (Hyndman & Athanasopoulos, 2018), we included Naive, Seasonal Naive, AutoETS, AutoARIMA (Hyndman et al., 2008) and AutoTheta (Assimakopoulos & Nikolopoulos, 2000). Additionally, we compared against several neural forecasting baselines, including WaveNet (Oord et al., 2016), DeepAR (Salinas et al., 2020), N-BEATS (Oreshkin et al., 2020), TFT (Lim et al., 2021), DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2023), N-HiTS (Challu et al., 2023), and GPT4TS (Zhou et al., 2023a). On Benchmark II (i.e., zero-shot datasets for Chronos models), we also evaluated against ForecastPFN (Dooley et al., 2023) which is a pretrained transformer model trained only on synthetic time series data.
@@ 5.5 Main Results
In this section, we present our main results on 42 datasets, which comprise Benchmark I (15 datasets) and Benchmark II (27 datasets). Chronos models surpass both classical statistical baselines and task-specific deep learning models on the in-domain datasets (Benchmark I; see Section 5.5.1). On the zero-shot datasets (Benchmark II; Section 5.5.2), Chronos models comfortably outperform statistical baselines, while performing on par with the best deep learning models trained on these tasks. With an inexpensive fine-tuning regimen, our Chronos-T5 (Small) model achieves the top spot on Benchmark II, significantly outperforming all baselines.
@@ 5.5.1 Benchmark I: In-domain Results
Figure 4: Performance of different models on Benchmark I, comprising 15 datasets also included in the training data of Chronos models. This benchmark showcases the in-domain performance of Chronos models against local statistical models, which fit parameters individually for each time series, and task-specific models that train a separate model for each task. The probabilistic (WQL) and point (MASE) forecasting metrics are normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS) are only compared based on MASE.
Benchmark I comprises 15 datasets that were also part of the training data of Chronos models, i.e., this benchmark evaluates the in-domain performance of Chronos models (see Table 2). Figure 4 summarizes the probabilistic and point forecasting performance for all models on the held-out test windows, in terms of their aggregated relative scores, computed as described in Section 5.4. The bigger Chronos-T5 models (Base and Large) significantly outperform baseline models, obtaining the best aggregated relative scores and average ranks (Figure 18 in Appendix E). These models not only perform better than local models (e.g., AutoETS and AutoARIMA), but they also perform better than task-specific deep learning models trained or fine-tuned for each dataset (e.g., PatchTST and DeepAR).
The smaller Chronos-T5 models (Mini and Small) and Chronos-GPT2 also perform better than the majority of baselines, with the exception of PatchTST. Task-specific deep learning models, trained across multiple time series for a specific task, perform better than local statistical models that fit parameters for each time series. Interestingly, the Seasonal Naive baseline performs competitively against other local models on this benchmark, suggesting that the datasets in this benchmark exhibit strong seasonal patterns. This is unsurprising since a majority of these datasets belong to domains such as energy and transport that tend to be highly seasonal in nature. The raw WQL and MASE values for individual datasets summarized in Figure 4 can be found in Tables 6 and 7 in Appendix E.
@@ 5.5.2 Benchmark II: Zero-shot Results
Figure 5: Performance of different models on Benchmark II, comprising 27 datasets not seen by Chronos models during training. This benchmark provides insights into the zero-shot performance of Chronos models against local statistical models, which fit parameters individually for each time series, task-specific models trained on each task, and the pretrained ForecastPFN model. The probabilistic (WQL) and point (MASE) forecasting metrics were normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS and ForecastPFN) are only compared based on MASE.
@@
Nevertheless, despite never having seen these datasets during training, Chronos models significantly outperform local statistical models. On probabilistic forecasting (aggregate relative WQL), Chronos models achieve the 2nd and 3rd spots, performing better than most task-specific models that have been trained on these tasks. Chronos-T5 (Large) places 3rd in terms of the point forecasting performance, narrowly losing the 2nd spot to N-HiTS. Chronos models also significantly outperform ForecastPFN, a recently proposed zero-shot forecaster, and even GPT4TS, which fine-tunes a pretrained GPT-2 model on each dataset. The raw WQL and MASE values for individual datasets summarized in Figure 5 can be found in Tables 8 and 9 in Appendix E.
@@ 5.6 Analysis of Hyperparameters
TSMix augmentations.
As described in Section 5.2, we trained Chronos models on TSMix augmentations rather than directly on the original time series. In this experiment, we investigate whether using TSMix augmentations is advantageous for downstream performance. Figure 10a compares the performance of Chronos-T5 (Small, 46M) models trained with and without TSMix augmentations. The model trained on TSMix augmentations obtains similar in-domain performance to the model trained without augmentations. However, the zero-shot performance improves when using TSMix augmentations. This suggests that TSMix enchances the diversity of training data which leads to improved performance on unseen datasets. Figure 10a also shows that the zero-shot performance obtains an additional boost with the inclusion of synthetic data. We investigate this further in the next experiment.
Figure 10: (a) Comparison of in-domain and zero-shot performance of Chronos-T5 (Small) models trained with and without TSMix augmentations. (b) In-domain and zero-shot performance of Chronos-T5 (Small) models with varying proportion of KernelSynth data in the training corpus.
@@
We systematically explored the impact of KernelSynth on downstream model performance. We trained Chronos-T5 (Small, 46M) models with time series sampled from TSMix augmentations and KernelSynth data in different ratios, ranging from 0% (i.e., trained solely on TSMix augmentations) to 100% synthetic data.
@@ 5.7 Qualitative Analysis and Limitations
We generated time series following linear and exponential trends: Chronos-T5 (Base) predicts the linear trend accurately but struggles with the exponential trend, as shown in Figure 12b. This may be due to a limited representation of exponential trends in the training data. A potential resolution for generating better forecasts for time series with exponential trends is to perform logarithmic scaling before feeding the time series into Chronos models. We also observed that Chronos models tend to underestimate the trend when the context is not sufficiently long. This phenomenon is depicted in Figure 13 where the model forecasts the pattern correctly but underpredicts the trend when a short context is provided. However, with a longer context, the model picks up the correct pattern and trend.
In our analysis, we observed that Chronos models recognize seasonal patterns in time series particularly well. We generated purely seasonal time series using sinusoids with different frequencies. As shown in Figure 12c, Chronos-T5 (Base) precisely forecasts both time series. When fundamental patterns such as trend and seasonality are combined, either additively or multiplicatively, Chronos forecasts them accurately. This is demonstrated in Figure 12d on time series generated via addition and multiplication of a linear function with a sinusoid.

### v2

@@ 1 Introduction
For the development of a useful general-purpose time series forecasting model, the scarcity of publicly available time series datasets, both in quantity and quality, is arguably more critical than the modeling framework. In addition to the comprehensive collection of public datasets we used to train Chronos, a central aspect of our approach is the integration of data augmentation strategies, including TSMixup and KernelSynth. TSMixup randomly samples a set of base time series from different training datasets, and generates new time series based on a convex combination of them; KernelSynth uses Gaussian processes to generate synthetic time series by randomly composing kernel functions. These techniques address the inherent limitations of small training datasets in time series forecasting, enhancing model robustness and generalization.
@@ 2 Background and Related Work
Time series forecasting can be addressed with a variety of different methods which can be broadly categorized into classical forecasting methods and deep learning methods. Classical forecasting methods such as ETS, ARIMA (Hyndman et al., 2008), Theta (Assimakopoulos & Nikolopoulos, 2000) fit a separate model to each time series independently (hence referred to as local models). In contrast, deep learning forecasting models learn across time series in a given dataset (and are called global models). These methods leverage advances in deep learning, such as RNNs which are used by DeepState (Rangapuram et al., 2018), DeepFactor  (Wang et al., 2019), DeepAR (Salinas et al., 2020), TimeGrad (Rasul et al., 2021), and transformers which are used by TFT (Lim et al., 2021) and PatchTST (Nie et al., 2023). Apart from the choice of architecture, these approaches differ in the way they model the target, with some modeling the density function while others directly predicting a set of quantiles (Wen et al., 2017; Gasthaus et al., 2019; Park et al., 2022). Nevertheless, not all models produce probabilistic forecasts: notably, models such as Informer (Zhou et al., 2021) and DLinear (Zeng et al., 2023) only produce point forecasts.
@@ 3.2 Objective Function
Opting for a categorical output distribution offers two key advantages. Firstly, it requires no modification to the language model architecture or training objective, enabling the use of popular language modeling libraries and the utilities they provide out of the box (Wolf et al., 2020). Secondly, it imposes no restrictions on the structure of the output distribution, allowing the model to learn arbitrary distributions, including multimodal ones. This flexibility proves especially valuable for a pretrained model, as time series datasets from diverse domains may follow distinct output distribution patterns.
@@ 4 Data Augmentation
4.1 TSMixup: Time Series Mixup
Figure 2: An illustration of TSMixup augmentation for $k=\{1,2,3\}$. TSMixup improves pattern diversity by taking weighted combinations of randomly-sampled time series from different datasets.
Mixup (Zhang et al., 2017) is a data augmentation scheme proposed in the context of image classification. It generates convex combinations of random image pairs and their labels from the training dataset, which alleviates issues such as memorization and overfitting in deep learning models. Existing works (Carmona et al., 2021; Zhou et al., 2023b) have extended Mixup to the time series domain. Building upon these works, we propose TSMixup, which generalizes the idea of Mixup to more than two datapoints. Concretely, TSMixup randomly samples $k\sim\mathcal{U}\{1,K\}$ time series of a specific length, $l\sim\mathcal{U}\{l_{\min},l_{\max}\}$, from the training datasets, scales them, and takes their convex combination,
$\tilde{{\bm{x}}}^{\mathrm{TSMixup}}_{1:l}=\sum_{i=1}^{k}\lambda_{i}\tilde{{\bm{x}}}^{(i)}_{1:l},$ (3)
where $\tilde{{\bm{x}}}^{(i)}_{1:l}$ denotes the $i$-th scaled time series. The combination weights, $[\lambda_{1},\dots,\lambda_{k}]$, are sampled from a symmetric Dirichlet distribution, $\mathrm{Dir}(\alpha)$. The complete pseudocode of TSMixup can be found in Algorithm 1 in Appendix A. Intuitively, TSMixup enhances the diversity of data by combining patterns from different time series. Figure 2 shows example augmentations generated by TSMixup and illustrates how different patterns are mixed.
@@ 4.2 KernelSynth: Synthetic Data Generation using Gaussian Processes
While TSMixup improves pattern diversity, it may still prove insufficient for training a generalist time series model, especially when real data is limited. To further supplement the training dataset, we propose KernelSynth, a method to generate synthetic time series using Gaussian processes (GPs). KernelSynth is inspired by the Automatic Statistician (Duvenaud et al., 2013), where a compositional search over a space of GP kernels is performed to explain the structure of a time series. We use the inverse of this process — randomly compose GP kernels to generate new time series.
@@ 5 Experiments
Pretraining-only 13 795,936 pretraining
Benchmark I 15 97,272 pretraining and in-domain evaluation Naive, SeasonalNaive, AutoETS, AutoTheta, AutoARIMA, DeepAR, TFT, PatchTST, DLinear, WaveNet, N-BEATS, N-HiTS, GPT4TS, Lag-Llama, Moirai-1.0-R
Benchmark II 27 190,674 zero-shot evaluation All the above, LLMTime and ForecastPFN
@@ 5.2 Training Corpus and Protocols
We trained T5 models of 4 sizes, [^1] Our code and model checkpoints are available at https://github.com/amazon-science/chronos-forecasting. namely, Mini (20M), Small (46M), Base (200M) and Large (710M), and the GPT-2 base model (90M), on 10M TSMixup augmentations (see Section 4.1) generated from the 28 training datasets, with $K=3$ in Algorithm 1, and 1M synthetic time series generated using Gaussian processes (see Section 4.2). Note that with this setup, original time series are adequately represented since they are included in the TSMixup augmentations with probability $1/3$. We sampled time series from the augmentations and synthetic data in the ratio 9:1 during training. Each model is trained with an effective batch size of 256 sequences, using distributed data parallelism and gradient accumulation, whenever necessary. These sequences are constructed by slicing random windows from the time series, and then scaling and quantizing them into equal-sized bins within the interval $[l{=}-15$, $r{=}15]$, as described in Section 3.1. The context length of the sequences was set to 512, the default for T5 models, and the prediction length is set to 64, a value greater than the prediction lengths of all tasks we consider in our evaluation.
@@ 5.3 Baselines
We assessed the performance of Chronos models against a variety of time series forecasting baselines. From statistical forecasting literature (Hyndman & Athanasopoulos, 2018), we included Naive, Seasonal Naive, AutoETS, AutoARIMA (Hyndman et al., 2008) and AutoTheta (Assimakopoulos & Nikolopoulos, 2000). Additionally, we compared against several neural forecasting baselines, including WaveNet (Oord et al., 2016), DeepAR (Salinas et al., 2020), N-BEATS (Oreshkin et al., 2020), TFT (Lim et al., 2021), DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2023), N-HiTS (Challu et al., 2023), and GPT4TS (Zhou et al., 2023a). Furthermore, from the recently proposed pretrained time series models, we included the ones with publicly available weights: Lag-Llama (Rasul et al., 2023) and Moirai-1.0-R (Woo et al., 2024). On Benchmark II (i.e., zero-shot datasets for Chronos models), we also evaluated against two zero-shot methods: ForecastPFN (Dooley et al., 2023) which is a transformer model pretrained only on synthetic time series data and LLMTime (Gruver et al., 2023) which uses LLMs for zero-shot forecasting.
@@ 5.5 Main Results
In this section, we present our main results on 42 datasets, which comprise Benchmark I (15 datasets) and Benchmark II (27 datasets). Chronos models surpass classical statistical baselines, task-specific deep learning models, and other pretrained models on the in-domain datasets (Benchmark I; see Section 5.5.1). On the zero-shot datasets (Benchmark II; Section 5.5.2), Chronos models comfortably outperform statistical baselines and other pretrained models, while performing on par with the best deep learning models trained on these tasks. With an inexpensive fine-tuning regimen, our Chronos-T5 (Small) model achieves the top spot on Benchmark II, significantly outperforming all baselines.
@@ 5.5.1 Benchmark I: In-domain Results
Figure 4: Performance of different models on Benchmark I, comprising 15 datasets also included in the training data of Chronos models. This benchmark showcases the in-domain performance of Chronos models against local statistical models, which fit parameters individually for each time series, task-specific models that train a separate model for each task, and pretrained models trained on a large corpus of time series data. Pretrained Models (Other) indicates that the in-domain setting does not apply to these models as some datasets in Benchmark I were not part of their training corpus and (or) they were trained on the test sets of some datasets in Benchmark I. The probabilistic (WQL) and point (MASE) forecasting metrics are normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS) are only compared based on MASE.
Benchmark I comprises 15 datasets that were also part of the training data of Chronos models, i.e., this benchmark evaluates the in-domain performance of Chronos models (see Table 2). Figure 4 summarizes the probabilistic and point forecasting performance for all models on the held-out test windows, in terms of their aggregated relative scores, computed as described in Section 5.4. The bigger Chronos-T5 models (Base and Large) significantly outperform baseline models, obtaining the best aggregated relative scores and average ranks (Figure 18 in Appendix E). These models not only perform better than local models (e.g., AutoETS and AutoARIMA), but they also perform better than task-specific deep learning models trained or fine-tuned for each dataset (e.g., PatchTST and DeepAR) and other pretrained models (e.g., Lag-Llama and Moirai-1.0-R).
The smaller Chronos-T5 models (Mini and Small) and Chronos-GPT2 also perform better than the majority of baselines, with the exception of PatchTST. Between the two baseline pretrained models studied in this experiment, Moirai-1.0-R clearly outperforms Lag-Llama. Notably, the best Moirai-1.0-R model (Large, 311M) is still outperformed by the smallest Chronos-T5 model (Mini, 20M) even though Moirai-1.0-R models were trained on a significantly larger corpus of time series data. Task-specific deep learning models, trained across multiple time series for a specific task, perform better than local statistical models that fit parameters for each time series. Interestingly, the Seasonal Naive baseline performs competitively against other local models on this benchmark, suggesting that the datasets in this benchmark exhibit strong seasonal patterns. This is unsurprising since a majority of these datasets belong to domains such as energy and transport that tend to be highly seasonal in nature. The raw WQL and MASE values for individual datasets summarized in Figure 4 can be found in Tables 6 and 7 in Appendix E.
@@ 5.5.2 Benchmark II: Zero-shot Results
Figure 5: Performance of different models on Benchmark II, comprising 27 datasets not seen by Chronos models during training. This benchmark provides insights into the zero-shot performance of Chronos models against local statistical models, which fit parameters individually for each time series, task-specific models trained on each task, and pretrained models trained on a large corpus of time series data. Pretrained Models (Other) indicates that the zero-shot setting does not apply to these models as they were trained on some datasets in Benchmark II. The probabilistic (WQL) and point (MASE) forecasting metrics were normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS and ForecastPFN) are only compared based on MASE.
@@
Nevertheless, despite never having seen these datasets during training, Chronos models significantly outperform local statistical models. On probabilistic forecasting (aggregate relative WQL), Chronos models achieve the 2nd and 3rd spots, performing better than most task-specific models that have been trained on these tasks. Chronos-T5 (Large) places 3rd in terms of the point forecasting performance, narrowly losing the 2nd spot to N-HiTS. Chronos models also significantly outperform other pretrained models such as Moirai-1.0-R, Lag-Llama, LLMTime, and ForecastPFN, and even GPT4TS, which fine-tunes a pretrained GPT-2 model on each dataset. Moirai-1.0-R obtains the best performance after Chronos although the evaluation setup may have been advantageous for Moirai-1.0-R as many datasets in Benchmark II were part of its pretraining corpus. The raw WQL and MASE values for individual datasets summarized in Figure 5 can be found in Tables 8 and 9 in Appendix E.
@@ 5.6 Analysis of Hyperparameters
TSMixup augmentations.
As described in Section 5.2, we trained Chronos models on TSMixup augmentations rather than directly on the original time series. In this experiment, we investigate whether using TSMixup augmentations is advantageous for downstream performance. Figure 10a compares the performance of Chronos-T5 (Small, 46M) models trained with and without TSMixup augmentations. The model trained on TSMixup augmentations obtains similar in-domain performance to the model trained without augmentations. However, the zero-shot performance improves when using TSMixup augmentations. This suggests that TSMixup enchances the diversity of training data which leads to improved performance on unseen datasets. Figure 10a also shows that the zero-shot performance obtains an additional boost with the inclusion of synthetic data. We investigate this further in the next experiment.
Figure 10: (a) Comparison of in-domain and zero-shot performance of Chronos-T5 (Small) models trained with and without TSMixup augmentations. (b) In-domain and zero-shot performance of Chronos-T5 (Small) models with varying proportion of KernelSynth data in the training corpus.
@@
We systematically explored the impact of KernelSynth on downstream model performance. We trained Chronos-T5 (Small, 46M) models with time series sampled from TSMixup augmentations and KernelSynth data in different ratios, ranging from 0% (i.e., trained solely on TSMixup augmentations) to 100% synthetic data.
@@ 5.7 Qualitative Analysis and Limitations
We generated time series following linear and exponential trends: Chronos-T5 (Base) predicts the linear trend accurately but struggles with the exponential trend, as shown in Figure 12b. This may be due to a limited representation of exponential trends in the training data. A potential resolution for generating better forecasts for time series with exponential trends is to perform logarithmic scaling before feeding the time series into Chronos models. We also observed that Chronos models tend to underestimate the trend when the context is not sufficiently long. This phenomenon is depicted in Figure 13 where the model forecasts the pattern correctly but underpredicts the trend when a short context is provided. However, with a longer context, the model picks up the correct pattern and trend. In our analysis, we observed that Chronos models recognize seasonal patterns in time series particularly well. We generated purely seasonal time series using sinusoids with different frequencies. As shown in Figure 12c, Chronos-T5 (Base) precisely forecasts both time series. When fundamental patterns such as trend and seasonality are combined, either additively or multiplicatively, Chronos forecasts them accurately. This is demonstrated in Figure 12d on time series generated via addition and multiplication of a linear function with a sinusoid.
(a) AR(1)
(b) AR(4)